# Quantum typicality approach to energy flow between two spin-chain domains at different temperatures

### v1

@@ V Results
Now, we turn to our numerical results, where we compare the energy current for various subsystem configurations to the value from CFT in Eq. (24). We first focus on the coupling of (i) two XX chains, (ii) two critical Ising chains, and (iii) one XX chain to a critical Ising chain. In CFT, the XX chain has the central charge $c=1$ while the transverse-field Ising chain at the critical point has the central charge $c=1/2$ [74].
@@
As one of our central results, we now calculate the steady-state energy current $J_{0}$ for a fixed initial temperature of $T_{L}/J=1/8$ of the left subsystem and different initial temperatures of the right subsystem, see Fig. 4. We compare our numerics to the CFT values in Eq. (24) for $c=1/2$ and $c=1$. For all considered subsystem configurations (i) – (iii), we have a convincing agreement between our numerical results and CFT. Also our data do not significantly depend on system size, as we show in Fig. 4 by also calculating the energy current for exemplary temperatures $T_{R}/J$ of the right subsystem for a total system size of $L=26$. Furthermore, for the coupling of an XX chain to a critical Ising chain, the energy current is determined by the system with the lower central charge, which is, in this scenario, the critical Ising chain, see Fig. 4 (c). Consequently, we observe a bottleneck effect [50], where the energy transport is limited by the subsystem with the fewer degrees of freedom.
@@
As the final subsystem configuration, we proceed to the (iv) coupling of two XXZ chains. We first focus on the case where all anisotropies are equal, i.e., $\Delta=\Delta_{L}=\Delta_{R}=\Delta_{C}$. We again benchmark our numerical calculation against existing results in the literature. In Fig. 5, we compare our numerics for a total system with $L=24$ sites, fixed temperatures of the left and right subsystem, and different anisotropies $\Delta$ to results from GHD. The numerical data for the site dependence of the local energy density (Fig. 5 (a)) as well as the data for the site dependence of the local energy current (Fig. 5 (b)) matches the results of Ref. [49] convincingly. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols). As shown in Fig. 5, the statistical error for a random state in the entire Hilbert space is slightly smaller than the statistical error for a random product state.
@@
Focusing again on the contact site, we calculate the energy current $J_{0}$ for the coupling of two XXZ chains with total system size $L=24$, fixed temperature $T_{L}/J=1/8$, and the anisotropies $\Delta=0.25,0.5$, and 1.01.01.0, see Fig. 6. For $-1\leq\Delta<1$, the ground state of the XXZ chain is gapless. The low-energy behavior of the XXZ chain in this regime is described by a CFT with central charge of $c=1$ [76]. Again, we recover the temperature dependence in Eq. (24) for all considered anisotropies, and our data agrees with the CFT result convincingly and also do not significantly depend on system size.
@@ VI Conclusion
To summarize, we discussed the method of dynamical quantum typicality, mostly used in linear response theory and at high temperatures, in less explored contexts. First, we applied it to energy flow between systems composed of two spin-chain domains at different temperatures and, second, to low temperatures. To test our method, we investigated the energy current in different spin-1/21/21 / 2 subsystem configurations, namely the coupling of (i) two XX chains, (ii) two critical Ising chains, (iii) one XX chain to a critical Ising chain, and (iv) two XXZ chains. We compared our numerical results to results from CFT and GHD. For all considered subsystem configurations and temperatures, we found a convincing agreement between our quantum typicality approach and these established methods. Therefore, we demonstrated the suitability of our approach for the study of bipartite systems at low temperatures. Further directions of research include additional models, such as the three-state quantum Potts chain, which in CFT has a central charge of $c=4/5$.

### v2

@@ V Results
Now, we turn to our numerical results, where we compare the energy current for various subsystem configurations to the value from CFT in Eq. (25). We first focus on the coupling of (i) two XX chains, (ii) two critical Ising chains, and (iii) one XX chain to a critical Ising chain. In CFT, the XX chain has the central charge $c=1$ while the transverse-field Ising chain at the critical point has the central charge $c=1/2$ Di Francesco et al. [1997].
@@
As one of our central results, we now calculate the steady-state energy current $J_{0}$ for a fixed initial temperature of $T_{L}/J=1/8$ of the left subsystem and different initial temperatures of the right subsystem, see Fig. 4. We compare our numerics to the CFT values in Eq. (25) for $c=1/2$ and $c=1$. For all considered subsystem configurations (i) – (iii), we have a convincing agreement between our numerical results and CFT. Also our data do not significantly depend on system size, as we show in Fig. 4 by also calculating the energy current for exemplary temperatures $T_{R}/J$ of the right subsystem for a total system size of $L=26$. Furthermore, for the coupling of an XX chain to a critical Ising chain, the energy current is determined by the system with the lower central charge, which is, in this scenario, the critical Ising chain, see Fig. 4 (c). Consequently, we observe a bottleneck effect Fischer et al. [2020], where the energy transport is limited by the subsystem with the fewer degrees of freedom.
Let us also give a technical remark here. It would be certainly interesting to explore even lower temperatures, e.g., to detect a region, where eventually DQT and CFT differ significantly from each other. However, in Fig. 4 we already consider a temperature $T_{R}/J$ = 0.05 or inverse temperature $\beta_{R}J=20$. To do so, we have to accurately approximate a small number of order $e^{-20}$ by an iterative procedure in imaginary time. An accurate approximation for even lower temperatures is quickly beyond the scope of such an interative procedure.
@@
As the final subsystem configuration, we proceed to the (iv) coupling of two XXZ chains. We first focus on the case where all anisotropies are equal, i.e., $\Delta=\Delta_{L}=\Delta_{R}=\Delta_{C}$. We again benchmark our numerical calculation against existing results in the literature. In Fig. 5, we compare our numerics for a total system with $L=24$ sites, fixed temperatures of the left and right subsystem, and different anisotropies $\Delta$ to results from GHD. The numerical data for the site dependence of the local energy density (Fig. 5 (a)) as well as the data for the site dependence of the local energy current (Fig. 5 (b)) matches the results of Ref. Bertini et al. [2016] convincingly. Here, the numerical data are shown for both, a random product state (open symbols) and a random state in the entire Hilbert space (closed symbols). As shown in Fig. 5, the statistical error for a random state in the entire Hilbert space is slightly smaller than the statistical error for a random product state. It would be also interesting to analyze states, which interpolate between these two states of no and maximum bipartite entanglement. But we leave this analysis for future research.
@@
Focusing again on the contact site, we calculate the energy current $J_{0}$ for the coupling of two XXZ chains with total system size $L=24$, fixed temperature $T_{L}/J=1/8$, and the anisotropies $\Delta=0.25,0.5$, and $1.0$, see Fig. 6. For $1\leq\Delta<1$, the ground state of the XXZ chain is gapless. The low-energy behavior of the XXZ chain in this regime is described by a CFT with central charge of $c=1$ Giamarchi [2003]. Again, we recover the temperature dependence in Eq. (25) for all considered anisotropies, and our data agrees with the CFT result convincingly and also do not significantly depend on system size.
@@ VI Conclusion
To summarize, we discussed the method of dynamical quantum typicality, mostly used in linear response theory and at high temperatures, in less explored contexts. First, we applied it to energy flow between systems composed of two spin-chain domains at different temperatures and, second, to low temperatures. To test our method, we investigated the energy current in different spin-$1/2$ subsystem configurations, namely the coupling of (i) two XX chains, (ii) two critical Ising chains, (iii) one XX chain to a critical Ising chain, and (iv) two XXZ chains. We compared our numerical results to results from CFT and GHD. For all considered subsystem configurations and temperatures, we found a convincing agreement between our quantum typicality approach and these established methods. Therefore, we demonstrated the suitability of our approach for the study of bipartite systems at low temperatures.
Interesting directions of future investigations include additional models, e.g., the three-state quantum Potts chain, which in CFT has a central charge of $c=4/5$. In addition, it is desirable to extend the investigation to gapped systems. Our expectation is that these systems are harder to treat by means of DQT, compared to the critial ones studied here.
@@ Acknowledgments
DATA AVAILABILITY
The data that support the findings of this article are openly available data .
Appendix A Details on the error bound
In the following, we provide details on the error bound for the variance, which can be found in, e.g., Knipschild [2021] in similar form.
Let us consider the variance
$\sigma^{2}=\overline{\langle\psi|O|\psi\rangle^{2}}-(\overline{\langle\psi|O|\psi\rangle})^{2}\;,$ (27)
where the overbar denotes the ensemble average over pure states $|\psi\rangle$. For a pure state $|\psi\rangle$ according to Eq. (17) one obtaines the variance
$\sigma^{2}$ $=$ $\sum_{i,j,k,l}\overline{c_{i}^{\ast}c_{j}c_{k}^{\ast}c_{l}}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{j}\rangle\langle m_{k}|\sqrt{\rho}O\sqrt{\rho}|m_{l}\rangle$ (28)
$-\sum_{i,k}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{i}\rangle\langle m_{k}|\sqrt{\rho}O\sqrt{\rho}|m_{k}\rangle\;.$
We now use the identity $\overline{c_{i}^{\ast}c_{j}c_{k}^{\ast}c_{l}}\approx\delta_{ij}\delta_{kl}+\delta_{il}\delta_{jk}$. Here, we can neglect the case $i=j=k=l$ for large dimensions $D\gg 1$ because this single case has only a contribution of order $\mathcal{O}(1/D)$. With this identity, we can simplify Eq. (28) to
$\sigma^{2}$ $\approx$ $\sum_{i,j}\langle m_{i}|\sqrt{\rho}O\sqrt{\rho}|m_{j}\rangle\langle m_{j}|\sqrt{\rho}O\sqrt{\rho}|m_{i}\rangle\;$ (29)
$=$ $\text{tr}[\rho O\rho O]\,.$
Next, we want to bound $\text{tr}[\rho O\rho O]$ from above. To this end, we use the spectral decomposition $p=\sum_{i}p_{i}|i\rangle\langle i|$ and get
$\text{tr}[\rho O\rho O]=\sum_{i,j}p_{i}p_{j}|\langle i|O|j\rangle|^{2}\;.$ (30)
Using
$(p_{i}-p_{j})^{2}=p_{i}^{2}-2p_{i}p_{j}+p_{j}^{2}\geq 0,$ (31)
we then obtain the inequality
$\text{tr}[\rho O\rho O]$ $\leq$ $\sum_{i,j}\frac{p_{i}^{2}+p_{j}^{2}}{2}|\langle i|O|j\rangle|^{2}$ (32)
$=$ $\sum_{i}p_{i}^{2}\langle i|O^{2}|i\rangle$
$=$ $\text{tr}[\rho^{2}O^{2}]\;.$
In the eigenbasis $|O_{j}\rangle$ of the operator $O$ we obtain for the trace on the r.h.s the estimation
$\text{tr}[\rho^{2}O^{2}]$ $=$ $\sum_{j}\langle O_{j}|\rho^{2}|O_{j}\rangle\langle O_{j}|O^{2}|O_{j}\rangle$ (33)
$\leq$ $O_{\text{max}}^{2}\,\text{tr}[\rho^{2}]\;,$
where $O_{\text{max}}=\text{max}[O_{j}]$ is the maximal eigenvalue. Thus, in summary, the variance can be bounded from above as
$\sigma^{2}\leq O_{\text{max}}^{2}\,\text{tr}[\rho^{2}]\,$, (34)
where $O_{\text{max}}$ of the local operator $O$ does not depend on system size $L$. Inserting for the density matrix $\rho$ the canonical ensemble leads to
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[(e^{-\beta H})^{2}]}{(\text{tr}[e^{-\beta H}])^{2}}$ (35)
or, equivalently,
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[(e^{-\beta(H-E_{0})})^{2}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}=\frac{\text{tr}[e^{-2\beta(H-E_{0})}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}$ (36)
with the ground-state energy $E_{0}$. Because
$\text{tr}[e^{-2\beta(H-E_{0})}]\leq\text{tr}[e^{-\beta(H-E_{0})}]\,$, (37)
one gets the expression
$\frac{\sigma^{2}}{O_{\text{max}}^{2}}\leq\frac{\text{tr}[e^{-\beta(H-E_{0})}]}{(\text{tr}[e^{-\beta(H-E_{0})}])^{2}}=\frac{1}{\text{tr}[e^{-\beta(H-E_{0})}]}\,.$ (38)
By introducing the effective dimension
$D_{\text{eff}}=\text{tr}[e^{-\beta(H-E_{0})}]\,$, (39)
this expression can be written as
$\sigma^{2}\leq O_{\text{max}}^{2}\,\frac{1}{D_{\text{eff}}}\,.$ (40)
Omitting the $L$-independent $O_{\text{max}}$ finally yields
$\sigma\leq b={\cal O}\left(\frac{1}{\sqrt{D_{\text{eff}}}}\right)\,$, (41)
cf., e.g., Jaklič and Prelovšek [2000]. Note that this bound can be derived analogously for a time-dependent operator $O(t)$ in the Heisenberg picture, because (i) $O(t)$ remains Hermitian under time evolution and (ii) the spectrum and thus $O_{\text{max}}$ is invariant under time evolution. Therefore, the bound in inquality (34) holds for $O(t)$ with the same constant $O_{\text{max}}$.